A chemist has one solution that is 80% acid and a second solution that is 30% acid. The chemist needs to mix some of both solutions to make 50 liters of a solution that is 62% acid. How much of each solution must he use?
Answer by jorel1380(3719)

50x.62=31
x+y=50
.8x+.3y=31
8x+3y=310
3x+3y=150
5x=160
x=32
y=18
The chemist needs 32 liters of 80% acid and 18 liters of 30% acid..